1.09g of a monobromoalkane, A that is immiscible with water was refluxed with excess aqueous sodium hydroxide. The liquid A was dissolved and the resulting solution was neutralized with nitric acid. Excess silver nitrate solution was added and a precipitate of silver bromide was formed. The precipitate was filtered, washed & dried. The precipitate weighed 1.99g.

The filtrate was distillled and a high boling point liquid B was obtained. b contained 52.2% carbon, 13% hydrogen and 34.8% oxygen by mass.

C = 12 H = 1 O = 16, Br = 80 Ag = 108

(a) Sketch a reflux apparatus

(b) Why was the solution not neutralised with hydrochloric acid?

(c) Write equations for the reaction taking place.


1
Expert's answer
2021-03-02T04:54:57-0500

A is bromoethane, B is ethanol:

MB = 16 / 0.348 = 46;

46 - 16 = 30 "\implies" C2H6

a)


b) Probably there should be nutric acid in the question. There was a solution of ethanol, excess NaOH and NaBr

c) C2H5Br + NaOH = C2H5OH + NaBr

NaOH + HNO3 = H2O + NaNO3

AgNO3 + NaBr = AgBr + NaNO3